User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Meeting 12th October

Chair: Tiege

Present: Tiege, Jamie, Lucia, Liz, Ellen, Mas, Patrick, Jesse

Apologies: Bryn, Jonny, Ross, Arwin, Lucas


General Issues:

  1. How is everyone going with Hobart 26m shifts currently, are people confused as to what back-ends should be being used or is that information getting conveyed properly?
    • Hopefully number of backends will be consolidated down to only two. MKV being retired.
    • Before starting any shift, check emails to make sure no specific information has been sent to you.
  2. Isn't LBA 2017OCT session soon?
    • No LBA observing due until next year!
  3. Cont17 roster - Nov 28 - Dec 12.
    • Ellen! Start the rostering!
    • 3 stations, two different schedules
    • Check roster as shift times may be changed.
    • 9am-3pm, 3pm-11pm and 11pm-9am
  4. Emails not getting sent through
    • Jamie has now fixed this issue, so the IVS ops email lists are correctly updated now.

Handover Issues:

  1. Ho: Vacuum pressure leak for receiver cooling system (AUA029)
    • Fixed! Caused by the weather warming.
  2. Ke: Mick changed disk in middle of recording, was the mystery disk that was originally being recorded to every identified? (R4811)
    • This was sorted, thanks to Warren.
    • Make sure to check module allocation sheet.
  3. Yg: Pointing offsets/tracking errors, Note from Arwin in handover notes: “During setup, noticed yg had large offsets in the antenna checks, had to the the pointing checks again and the offsets were better the second time around. This has happened a few times, should this be raised at the meeting?”
    • The pointing should be OK, Jamie is going to have a squiz.
/home/www/auscope/opswiki/data/pages/operations/agenda2017_10_12.txt · Last modified: 2017/10/12 02:25 by Tiege McCarthy